# QCD axion dark matter in the dark dimension

### v1

@@
Abstract
The recently proposed dark dimension scenario reveals that the axions can be localized on the Standard Model brane, thereby predicting the quantum chromodynamics (QCD) axion decay constant from the weak gravity conjecture: $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ is the five-dimensional Planck mass. When combined with observational lower bounds, this implies that $f_{a}$ falls within a narrow range $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, corresponding to the axion mass $m_{a}\sim 10^{-3}-10^{-2}\,\rm eV$. At this scale, the QCD axion constitutes a minor fraction of the total cold dark matter (DM) density $\sim 10^{-3}-10^{-2}$. In this work, we investigate the issue of QCD axion DM within the context of the dark dimension and demonstrate that the QCD axion in this scenario can account for the entire DM abundance through a simple two-axion mixing mechanism. Here we consider the resonant conversion of an axion-like particle (ALP) into a QCD axion. We find that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the QCD axion in the dark dimension scenario can account for the overall DM.
Keywords:
Axions and ALPs, Particle Nature of Dark Matter, Phase Transitions in the Early Universe
preprint: ITP-24-277arxiv: 2412.xxxx
@@ 1 Introduction
The axion is a pseudoscalar field that was initially foreseen by the Peccei-Quinn (PQ) mechanism Peccei:1977hh ; Peccei:1977ur . This mechanism is characterized by a spontaneously broken global $\rm U(1)_{PQ}$ symmetry, wherein the axion, denoted as $\theta$, undergoes a transformation of $\theta\to\theta+c$, with $c$ being a constant. The primary objective of this mechanism is to address the strong CP problem in the Standard Model (SM) in a dynamic manner; consequently, it is also referred to as the quantum chromodynamics (QCD) axion Weinberg:1977ma ; Wilczek:1977pj ; Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq . This axion acquires an exceedingly small mass from the QCD non-perturbative effects tHooft:1976rip ; tHooft:1976snw . When the PQ symmetry is broken by the QCD instanton, it can lead to the induction of an axion potential. As a result, the QCD axion can stabilize at the CP conservation minimum, thereby dynamically resolving the strong CP problem Hook:2018dlk .
@@
where $f_{a}$ denotes the QCD axion decay constant. In the absence of fine-tuning of the initial misalignment angle, a large $f_{a}$ would lead to an overproduction of QCD axion DM, thereby establishing this upper bound. On the other hand, the lower bound of $f_{a}$ is approximately $\sim 10^{8}-10^{9}\,{\rm GeV}$, which stems from astrophysical observation constraints, including the duration of the neutrino burst from supernova SN 1987A Raffelt:1987yt ; Turner:1987by ; Mayle:1987as and the cooling rate of neutron stars Leinson:2014ioa ; Hamaguchi:2018oqw ; Buschmann:2021juv , among others.
@@
In general, the axions originating from four-dimensional models can be referred to as the “type I axions”. Additionally, the “type II axions”, which encompass numerous axion-like particles (ALPs) and also QCD axions, can arise from higher-dimensional gauge fields Witten:1984dg ; Green:1984sg ; Choi:2003wr . In the context of string theory Svrcek:2006yi ; Conlon:2006tq ; Cicoli:2012sz , lower-dimensional axions can emerge by integrating higher-dimensional gauge fields over cycles within the compactified space. The resultant axion decay constant is dictated by the internal geometry of the compactification, often referred to as the model-dependent axion. Furthermore, there exists the model-independent axion, which is unaffected by the structure of the internal manifold. See also $\rm e.g.$ refs. Reece:2024wrn ; Choi:2024ome for recent discussion on axions in the extra dimension.
@@
The recently proposed dark dimension scenario Montero:2022prj predicts a single large extra dimension with the range $L_{5}\sim 1-10\,\rm\mu m$, leading to a specific corner of the quantum gravity landscape that corresponds to an asymptotic region of the field space. This scenario is motivated by the smallness of dark energy, informed by Swampland principles Vafa:2005ui ; Ooguri:2006in , and constrained by observational data. In this context, the SM is localized on a codimension-one brane within the five-dimensional spacetime. Furthermore, this scenario gives rise to intriguing phenomenology related to the dark dimension, as explored in various studies Gonzalo:2022jac ; Law-Smith:2023czn ; Obied:2023clp ; Anchordoqui:2022svl ; Anchordoqui:2022tgp ; Anchordoqui:2022txe ; Anchordoqui:2023etp ; Anchordoqui:2023tln ; Anchordoqui:2024akj ; Anchordoqui:2024dxu ; Anchordoqui:2024tdj . See also ref. Vafa:2024fpx for a recent review.
@@
The axions in the dark dimension have recently been investigated in ref. Gendler:2024gdo . Within this context, it is a natural consideration to localize the QCD axion on the SM brane. By applying the weak gravity conjecture (WGC) Arkani-Hamed:2006emk to the QCD axion, we can obtain an upper bound for the axion decay constant, $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ represents the five-dimensional Planck mass. Additionally, one can derive this inequality by considering axion propagation throughout the entire five-dimensional bulk. On the observational front, constraints indicate a lower limit for the classical QCD axion window, specifically $f_{a}\gtrsim 10^{8}-10^{9}\,{\rm GeV}$. Consequently, if the dark dimension scenario accurately describes our Universe, the QCD axion localized on the SM brane must possess a narrow range of axion decay constant Gendler:2024gdo . This range is expressed as follows
@@
Notably, eq. (2) falls precisely within the classical QCD axion window. By employing the misalignment mechanism, it can be determined that the QCD axion within this range comprises a small fraction of the overall DM density, ranging from $\sim 10^{-3}$ to $10^{-2}$.
In this work, we investigate the issue of QCD axion DM within the dark dimension. We demonstrate that the QCD axion in the dark dimension can fully account for the abundance of DM, which is achieved through a straightforward two-axion mixing mechanism. The concept of axiverse, a Universe populated by a multitude of axions, including both QCD axions and ALPs, has been discussed extensively in the literature Arvanitaki:2009fg ; Demirtas:2021gsq . Within this framework, it is plausible to consider the cosmological evolution of multiple axions in the early Universe. Here, we focus on the resonant conversion of ALP into QCD axion within the context of the dark dimension, a process that can take place prior to the critical temperature of the QCD phase transition. This phenomenon of two-axion mixing has been extensively studied over the past decade, and the resulting QCD axion DM abundance can be either suppressed or enhanced Hill:1988bu ; Cyncynates:2023esj ; Kitajima:2014xla ; Daido:2015cba ; Ho:2018qur ; Li:2023xkn ; Li:2023uvt ; Li:2024okl ; Murai:2024nsp . Our findings indicate that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the resulting QCD axion from the resonant conversion can indeed constitute the entirety of the cold DM abundance. Finally, we brief discuss the adiabatic condition in the axion resonant conversion.
The rest of this paper is structured as follows. In section 2, we provide a concise overview of the QCD axion within the context of the dark dimension. In section 3, we discuss the QCD axion DM in this scenario and demonstrate how the abundance of QCD axion DM is enhanced through the resonant conversion from ALP to QCD axion. Finally, the conclusion is given in section 4.
@@ 2.1 Axions from higher-dimensional gauge fields
Here we briefly discuss axions that originate from higher-dimensional gauge fields. For a more comprehensive understanding, please refer to recent refs. Reece:2024wrn ; Choi:2024ome .
Firstly, we demonstrate that the theory of an higher-dimensional $p$-form gauge field can generate massless four-dimensional axions. Assume that spacetime consists of $d=(4+n)$ dimensions, which is manifested as a warped product compactification
@@
where $b_{p}(Y)$ is the Betti number, and $\omega_{p}^{(i)}$ represents the corresponding cohomology class. In $d=(4+n)$ dimensions, the standard kinetic term for the $p$-form gauge field can be described by
@@
The diagonal entries of the kinetic matrix $k_{ij}$ can be interpreted as the squares of the axion decay constants, denoted by
@@
In the context of large gauge transformations, the redundancy manifests as a precise periodicity constraint on the four-dimensional axions, expressed as
@@
Above is the depiction of massless four-dimensional axions originating from higher-dimensional $p$-form gauge fields. Additionally, there are effects that contribute to a small mass for these four-dimensional axions, such as couplings to objects carrying a $A_{p}$ gauge charge, and Chern-Simons interactions involving $A_{p}$ with itself or other gauge fields. The existence of such axions represents a topological characteristic of the theory. Specifically, there is no requirement for supplementary frameworks such as supersymmetry to elucidate why an axion in an extra dimension is light; it is inherently exponentially suppressed in mass, as long as it does not acquire a tree-level mass contribution from the topological terms. Consequently, they alleviate the axion quality problem significantly Reece:2024wrn .
@@ 2.2 Axions in the dark dimension
In this subsection, we briefly review the QCD axion in the dark dimension.
The dark dimension scenario Montero:2022prj predicts one extra mesoscopic dimension in the micron range, which is achieved by applying various Swampland principles to the dark energy. The dark energy in Planck units $\Lambda$ is a hierarchically small parameter. Since the emergence of any small parameters within the framework of quantum gravity necessitates the presence of towers of light states, which can arise due to either some dimensions expanding to a large size or a critical string becoming light Lee:2019wij , the sole experimentally allowed model featuring large extra dimensions in this context is one that involves a single large extra dimension with the length
@@
The WGC Arkani-Hamed:2006emk asserts that, for compatibility with quantum gravity, there must exist a U(1) charged particle whose gravitational interaction is weaker than its U(1) gauge interaction. In Planck units, this translates to the existence of a “superextremal” particle that satisfies the following inequality
@@
The WGC extends its applicability beyond mere particles, encompassing any charged state associated with a $p$-form gauge field. Specifically, for any given $p$-form gauge field, there ought to exist a ($p-1$)-dimensional state, characterized by a charge-to-mass ratio that surpasses or equals that of an extremal ($p-1$) black brane. When applied to the 00-form gauge field, namely the axion, we have the inequality
@@
where $S_{\rm inst}$ represents the instanton action, and $f_{a}$ is the axion decay constant. It’s worth noting that, due to the absence of an extremal solution for the instanton, this inequality encompasses an undefined factor of order one. Given that the dilute instanton gas approximation necessitates $S_{\rm inst}\gtrsim 1$ to maintain control over the instanton expansion, we can deduce the axion WGC bound
@@
Now, let’s consider a four-dimensional brane localized in the fifth dimension and apply the WGC to particles localized on this brane. This yields
@@
In this context, it’s reasonable to consider the QCD axion as the DM candidate, which is the focus of this work. However, as discussed in Gendler:2024gdo , the QCD axion within this range can only contribute a small fraction to the overall DM density.
@@ 3.1 QCD axion dark matter
Here we discuss the QCD axion DM in the dark dimension scenario through the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah .
Since the extra-dimensional axion is not associated with the cosmological PQ phase transition, as discussed in ref. Reece:2023czb , we can consider it within the canonical pre-inflationary scenario where the PQ symmetry is spontaneously broken during inflation. The low-energy effective Lagrangian of the QCD axion, stemming from the QCD non-perturbative effects, can be described by
@@
where $T_{i,a}$ is the oscillation temperature
@@
$N_{a}\equiv\dfrac{\rho_{a}}{m_{a}}a^{3}\,,$ (34)
where $T_{0}$ is the current cosmic microwave background (CMB) temperature, and $a$ is the scale factor. The QCD axion energy density at $T_{0}$ is given by
$\rho_{a,0}=\frac{1}{2}m_{a,0}m_{a,i}f_{a}^{2}\left\langle\theta_{i}^{2}f(\theta_{i})\right\rangle\chi\left(\frac{a_{i,a}}{a_{0}}\right)^{3}\,,$ (35)
where $a_{x}$ denotes the scale factor at the temperature $T_{x}$. Then the current QCD axion DM abundance can be described by
@@
where $h\simeq 0.68$ is the reduced Hubble constant, $\rho_{\rm crit}$ is the critical energy density, $g_{*}(T)$ and $g_{*s}(T)$ are the numbers of effective degrees of freedom for the energy density and the entropy density, respectively. In order to explain the observed DM abundance, $\Omega_{\rm DM}h^{2}\simeq 0.12$ Planck:2018vyg , we require an $\sim\mathcal{O}(1)$ initial misalignment angle
@@
Now considering the QCD axion DM in the dark dimension scenario with a narrow axion decay constant $f_{a}$ ranging from $10^{9}$ to $10^{10}\,\rm GeV$, we can obtain a small fraction of the overall DM density Gendler:2024gdo
@@
Then, the situation begins to take on a captivating turn. This is because, in general, the issue we face is the overproduction of the QCD axion DM abundance. Specifically, when the decay constant of the QCD axion exceeds $10^{12}\,\rm GeV$ and we only consider an initial misalignment angle of order unity, the axion DM abundance will be overproduced. Therefore, in this scenario, fine-tuning the initial misalignment angle emerges as a viable resolution. However, as analyzed above, when considering the QCD axion in the context of dark dimension, the values of the axion decay constant within the range of $\sim 10^{8}-10^{9}\,{\rm GeV}$ are inadequate to explain the observed DM abundance. Therefore, our subsequent endeavor will be to explore how to enhance the QCD axion DM abundance for relatively smaller decay constants $f_{a}\sim 10^{8}-10^{9}\,{\rm GeV}$. Specifically, we will consider the effect of two-axion mass mixing to address this issue Cyncynates:2023esj ; Li:2024okl .

### v2

# ITP-CAS-24-277

@@
ITP-CAS-24-277
QCD Axion Dark Matter in the Dark
Dimension
Hai-Jun Li
Key Laboratory of Theoretical Physics, Institute of Theoretical Physics, Chinese Academy of Sciences, Beijing 100190, China
E-mail: lihaijun@itp.ac.cn
The recently proposed dark dimension scenario reveals that axions can be localized on the Standard Model brane, thereby predicting the quantum chromodynamics (QCD) axion decay constant from the Weak Gravity Conjecture: $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ is the five-dimensional Planck mass. When combined with observational lower bounds, this implies that $f_{a}$ falls within a narrow range $f_{a}\sim 10^{9}-10^{10}\,\rm GeV$, corresponding to the axion mass $m_{a}\sim 10^{-3}-10^{-2}\,\rm eV$. At this scale, the QCD axion constitutes a minor fraction of the total cold dark matter (DM) density $\sim 10^{-3}-10^{-2}$. In this work, we investigate the issue of QCD axion DM within the context of the dark dimension and demonstrate that the QCD axion in this scenario can account for the entire DM abundance through a simple two-axion mixing mechanism. Specifically, we consider the resonant conversion of an axion-like particle (ALP) into the QCD axion. We find that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the QCD axion in the dark dimension can account for the overall DM. The ALP required within this specific range may originate from the grand unification of gauge forces in the dark dimension.
April 14, 2025
@@ 1 Introduction
The axion is a pseudoscalar field that was initially foreseen by the Peccei-Quinn (PQ) mechanism Peccei:1977hh ; Peccei:1977ur . This mechanism is characterized by a spontaneously broken global $\rm U(1)_{PQ}$ symmetry, wherein the axion, denoted as $\theta$, undergoes a transformation of $\theta\to\theta+c$, with $c$ being a constant. The primary objective of this mechanism is to address the strong CP problem in the Standard Model (SM) in a dynamic manner; consequently, it is also referred to as the quantum chromodynamics (QCD) axion Weinberg:1977ma ; Wilczek:1977pj ; Kim:1979if ; Shifman:1979if ; Dine:1981rt ; Zhitnitsky:1980tq . This axion acquires a small mass from QCD non-perturbative effects tHooft:1976rip ; tHooft:1976snw . When the PQ symmetry is broken by the QCD instanton, it can lead to the induction of an axion potential. As a result, the QCD axion can stabilize at the CP conservation minimum, thereby dynamically resolving the strong CP problem Hook:2018dlk .
@@
where $f_{a}$ denotes the QCD axion decay constant. In the absence of fine-tuning of the initial misalignment angle, a large $f_{a}$ would lead to an overproduction of the QCD axion DM, thereby establishing this upper bound. On the other hand, the lower bound of $f_{a}$ is approximately $\sim 10^{8}-10^{9}\,{\rm GeV}$, which stems from astrophysical observation constraints, including the duration of the neutrino burst from supernova SN 1987A Raffelt:1987yt ; Turner:1987by ; Mayle:1987as and the cooling rate of neutron stars Leinson:2014ioa ; Hamaguchi:2018oqw ; Buschmann:2021juv , among others.
@@
In general, axions originating from four-dimensional models can be categorically designated as “type I axions."""" Additionally, “type II axions,"""" which encompass numerous axion-like particles (ALPs) as well as the QCD axion, can arise from higher-dimensional gauge fields Witten:1984dg ; Green:1984sg ; Choi:2003wr . In the context of string theory Svrcek:2006yi ; Conlon:2006tq , lower-dimensional axions can emerge by integrating higher-dimensional gauge fields over cycles within the compactified space. The resultant axion decay constant is dictated by the internal geometry of the compactification, often referred to as the model-dependent axion. Furthermore, there exists the model-independent axion, which remains steadfastly unaffected by the structure of the internal manifold. See also $\rm e.g.$ refs. Reece:2024wrn ; Choi:2024ome for recent discussions on axions in extra dimensions.
@@
The recently proposed dark dimension scenario Montero:2022prj predicts a single large extra dimension with the range $L_{5}\sim 1-10\,\rm\mu m$, leading to a specific corner of the quantum gravity landscape that corresponds to an asymptotic region of the field space. This scenario is motivated by the smallness of dark energy, informed by Swampland principles Vafa:2005ui ; Ooguri:2006in , and constrained by observational data. In this context, the SM is localized on a codimension-one brane within the five-dimensional spacetime. Furthermore, this scenario gives rise to intriguing phenomenology related to the dark dimension, as explored in various studies Gonzalo:2022jac ; Law-Smith:2023czn ; Obied:2023clp ; Anchordoqui:2022svl ; Anchordoqui:2022tgp ; Anchordoqui:2022txe ; Anchordoqui:2023etp ; Anchordoqui:2023tln ; Anchordoqui:2024akj ; Anchordoqui:2024dxu ; Anchordoqui:2024tdj ; BitaghsirFadafan:2023hwg ; Schwarz:2024tet ; Heckman:2024trz . See also ref. Vafa:2024fpx for a recent review.
@@
Axions in the dark dimension have recently been investigated in ref. Gendler:2024gdo . Within this context, it is natural to localize the QCD axion on the SM brane. By applying the Weak Gravity Conjecture (WGC) Arkani-Hamed:2006emk to the QCD axion, we can obtain an upper bound for the axion decay constant, $f_{a}\lesssim M_{5}\sim 10^{9}-10^{10}\,\rm GeV$, where $M_{5}$ represents the five-dimensional Planck mass. Additionally, one can derive this inequality by considering axion propagation throughout the entire five-dimensional bulk. On the observational front, constraints indicate a lower limit for the classical QCD axion window, specifically $f_{a}\gtrsim 10^{8}-10^{9}\,{\rm GeV}$. Consequently, if the dark dimension scenario accurately describes our Universe, the QCD axion localized on the SM brane must possess a narrow range of axion decay constant Gendler:2024gdo . This range is expressed as follows
@@
Notably, eq. (2) falls precisely within the established range of the classical QCD axion window. By employing the misalignment mechanism, it can be determined that the QCD axion within this specific range comprises a small fraction of the overall DM density, ranging from $\sim 10^{-3}$ to $10^{-2}$.
In this work, we investigate the issue of QCD axion DM within the context of the dark dimension. We demonstrate that the QCD axion in the dark dimension can fully account for the abundance of DM, which is achieved through a straightforward two-axion mixing mechanism. The concept of axiverse, a Universe populated by a multitude of axions, including both the QCD axion and a large number of ALPs, has been discussed extensively in the literature Arvanitaki:2009fg ; Acharya:2010zx ; Cicoli:2012sz ; Demirtas:2018akl ; Reig:2021ipa ; Demirtas:2021gsq ; Alexander:2024nvi . Within this framework, it is plausible to consider the cosmological evolution of multiple axions in the early Universe. Here, we focus on the resonant conversion of ALP into QCD axion in the dark dimension scenario, a process that can take place prior to the critical temperature of the QCD phase transition. This similar phenomenon of two-axion mixing has been extensively studied over the past decade, and the resulting QCD axion DM abundance can be either effectively suppressed or enhanced Hill:1988bu ; Cyncynates:2023esj ; Li:2024okl ; Li:2024kdy ; Kitajima:2014xla ; Daido:2015cba ; Ho:2018qur ; Li:2023det ; Li:2023xkn ; Li:2023uvt ; Murai:2024nsp . Our findings indicate that, in a scenario where the ALP possesses a mass of approximately $m_{A}\sim 10^{-5}\,\rm eV$ and a decay constant of $f_{A}\sim 10^{11}\,\rm GeV$, the resulting QCD axion from the resonant conversion can indeed constitute the entirety of the cold DM abundance. Lastly, we briefly discuss the adiabatic condition in axion resonant conversion and the possible origins of the ALP with a higher-scale decay constant.
The rest of this paper is structured as follows. In section 2, we provide a concise overview of the QCD axion within the context of the dark dimension. In section 3, we discuss the QCD axion DM in this scenario and demonstrate how the abundance of QCD axion DM is enhanced through the resonant conversion of ALP into QCD axion. Finally, the conclusion is presented in section 4.
@@ 2.1 Axions from higher-dimensional gauge fields
Here we introduce axions that originate from higher-dimensional gauge fields. For a more comprehensive understanding, please refer to recent refs. Reece:2024wrn ; Choi:2024ome .
Firstly, we demonstrate that the theory of a higher-dimensional $p$-form gauge field can generate massless four-dimensional axions. Assume that spacetime consists of $d=(4+n)$ dimensions, which is manifested as a warped product compactification
@@
where $b_{p}(Y)$ represents the Betti number, and $\omega_{p}^{(i)}$ represents the corresponding cohomology class. In $d=(4+n)$ dimensions, the standard kinetic term for the $p$-form gauge field can be described by
@@
The diagonal entries of the kinetic matrix $k_{ij}$ can be interpreted as the squares of axion decay constants, denoted by
@@
In the context of large gauge transformations, the redundancy manifests as a precise periodicity constraint on four-dimensional axions, expressed as
@@
Above is the depiction of massless four-dimensional axions originating from higher-dimensional $p$-form gauge fields. Additionally, there are various effects that contribute to a small mass for these four-dimensional axions, such as their couplings to objects carrying an $A_{p}$ gauge charge, as well as Chern-Simons (CS) interactions involving $A_{p}$ with itself or other gauge fields. The existence of such axions represents a topological characteristic inherent to the theory. Specifically, there is no requirement for supplementary frameworks such as supersymmetry to elucidate why an axion in an extra dimension is light; it is inherently exponentially suppressed in mass, as long as it does not acquire a tree-level mass contribution from topological terms. Consequently, these axions alleviate the axion quality problem significantly Reece:2024wrn .
@@ 2.2 Axions in the dark dimension
In this subsection, we briefly review the QCD axion in the dark dimension. The dark dimension scenario Montero:2022prj predicts one extra mesoscopic dimension in the micron range, which is achieved by applying various Swampland principles to the dark energy. The dark energy in Planck units $\Lambda$ is a hierarchically small parameter. Since the emergence of any small parameters within the framework of quantum gravity necessitates the presence of towers of light states, which can arise due to either some dimensions expanding to a large size or a critical string becoming light Lee:2019wij , the sole experimentally allowed model featuring large extra dimensions in this context is one that involves a single large extra dimension with the length
@@
The WGC Arkani-Hamed:2006emk asserts that, for compatibility with quantum gravity, there must exist a U(1) charged particle whose gravitational interaction is weaker than its U(1) gauge interaction. In Planck units, this translates to the existence of a “superextremal"""" particle that satisfies the following inequality
@@
The applicability of the WGC extends beyond mere particles, encompassing any charged state associated with a $p$-form gauge field. Specifically, for any given $p$-form gauge field, there ought to exist a ($p-1$)-dimensional state, characterized by a charge-to-mass ratio that surpasses or equals that of an extremal ($p-1$) black brane. When applied to the 00-form gauge field, namely the axion, we have the inequality
@@
where $S_{\rm inst}$ represents the instanton action, and $f_{a}$ is the axion decay constant. It is worth noting that, due to the absence of an extremal solution for the instanton, this inequality encompasses an undefined factor of order one. Given that the dilute instanton gas approximation necessitates $S_{\rm inst}\gtrsim 1$ to maintain control over the instanton expansion, we can obtain the axion WGC bound
@@
Now, let us consider a four-dimensional brane localized in the fifth dimension and apply the WGC to particles localized on this brane. This yields
@@
In this context, it is reasonable to consider the QCD axion as the DM candidate, which is the focus of this work. However, as discussed in Gendler:2024gdo , the QCD axion within this range can only contribute a small fraction to the overall DM density.
@@ 3.1 QCD axion dark matter
Here we discuss the QCD axion DM in the dark dimension scenario through the misalignment mechanism Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . Since the extra-dimensional axion is not associated with the cosmological PQ phase transition, as discussed in ref. Reece:2023czb , we can consider it within the canonical pre-inflationary scenario where the PQ symmetry is spontaneously broken during inflation. The low-energy effective Lagrangian of the QCD axion, stemming from QCD non-perturbative effects, can be described by
@@
where $T_{i,a}$ is the oscillation temperature Li:2023xkn
@@
$N_{a}\equiv\dfrac{\rho_{a}}{m_{a}}\mathfrak{a}^{3}\,,$ (34)
where $T_{0}$ is the current cosmic microwave background (CMB) temperature, and $\mathfrak{a}$ is the scale factor. The axion energy density at $T_{0}$ is given by
$\rho_{a,0}=\frac{1}{2}m_{a,0}m_{a,i}f_{a}^{2}\left\langle\theta_{i}^{2}f(\theta_{i})\right\rangle\chi\left(\frac{\mathfrak{a}_{i,a}}{\mathfrak{a}_{0}}\right)^{3}\,,$ (35)
where $\mathfrak{a}_{x}$ denotes the scale factor at the temperature $T_{x}$. Then the current QCD axion DM abundance can be described by Li:2023xkn
@@
where $h\simeq 0.68$ is the reduced Hubble constant, $\rho_{\rm crit}$ is the critical energy density, $g_{*}(T)$ and $g_{*s}(T)$ are the numbers of effective degrees of freedom for the energy density and the entropy density, respectively. In order to explain the observed DM abundance, $\Omega_{\rm DM}h^{2}\simeq 0.12$ Planck:2018vyg , we require an $\sim\mathcal{O}(1)$ initial misalignment angle Li:2023xkn
@@
Now, considering the QCD axion DM in the dark dimension scenario with a narrow axion decay constant $f_{a}$ ranging from $10^{9}$ to $10^{10}\,\rm GeV$, we can obtain a small fraction of the overall DM density Gendler:2024gdo
@@
Then, the situation starts to take a captivating turn. This is because, in general, the issue we face is the overproduction of the QCD axion DM abundance. Specifically, when the decay constant of the QCD axion exceeds $10^{12}\,\rm GeV$ and we only consider an initial misalignment angle of order unity, the axion DM abundance will be overproduced. Therefore, in this scenario, fine-tuning the initial misalignment angle emerges as a viable resolution. However, as analyzed above, when considering the QCD axion in the dark dimension scenario, the value of the axion decay constant within the range of $\sim 10^{9}-10^{10}\,{\rm GeV}$ is inadequate to explain the observed DM abundance. Therefore, our subsequent endeavor will be to explore how to enhance the QCD axion DM abundance for a relatively smaller decay constant $f_{a}\sim 10^{9}-10^{10}\,{\rm GeV}$. Specifically, we will consider the effect of two-axion mass mixing to address this issue Cyncynates:2023esj ; Li:2024okl .